# SURT 2.0: Advances in Transducer-based Multi-talker Speech Recognition

### v1

@@ I Introduction
For the case of streaming multi-talker ASR, two conceptually similar frameworks — streaming unmixing and recognition transducer (SURT) [20] and multi-turn RNNT (MT-RNNT) [21] — have been proposed that seek to extend neural transducers [22] for the task. This is beneficial since neural transducers (using RNNs or transformers) have become the standard modeling technique for on-device speech recognition [23, 24, 25]. Both these frameworks separate overlapping speech into multiple simultaneous branches (or channels), each of which is transcribed by a shared transducer. These models have been extended to handle long-form multi-turn recordings [26, 27], and to jointly perform speaker identification [28], endpointing [29], and segmentation [30]. In this paper, we refer to these class of models as SURT, but the same ideas should also be applicable to MT-RNNT.
@@ III-A Speech recognition with neural transducers
In conventional single-talker ASR, audio features for a segmented utterance $\mathbf{X}\in\mathbb{R}^{T\times F}$ are provided as input to the system, and we are required to predict the transcript $\mathbf{y}=(y_{1},\ldots,y_{U})$, where $y_{u}\in\Sigma$ denotes output units such as graphemes or word-pieces. For the case of discriminative training, this requires computing the conditional likelihood $P(\mathbf{y}|\mathbf{X})$ (or its log for numerical stability). For inference, we search for $\hat{\mathbf{y}}=\text{arg}\max_{\mathbf{y}}P(\mathbf{y}|\mathbf{X})$, often in a constrained search space using greedy or beam search. Transducers achieve this by marginalizing over the set of all alignments $\mathbf{a}\in\bar{\Sigma}^{T+U}$, where $\bar{\Sigma}=\Sigma\cup\{\phi\}$ and $\phi$ is called the blank label. Formally,
@@
Transducers parameterize $P(\mathbf{a}|\mathbf{X})$ with an encoder, a prediction network, and a joiner (see “recognition” component in Fig. 2a). The encoder maps $\mathbf{x}$ into hidden representations $\mathbf{f}_{1}^{T}$, while the prediction network maps $\mathbf{y}$ into $\mathbf{g}_{1}^{U}$. The joiner combines the outputs from the encoder and the prediction network to compute logits $\mathbf{z}_{t,u}$ which are fed to a softmax function to produce a posterior distribution over $\bar{\Sigma}$. Under assumptions of full context encoder, we can expand (1) as
@@
where $u(t)$ denotes the index in the label sequence at time $t$. We will denote the log of this expression as $\mathcal{L}_{\text{rnnt}}(\mathbf{y},\mathbf{z})$ (or simply $\mathcal{L}_{\text{rnnt}}$) for the remainder of this paper.
@@ III-B Multi-talker ASR with SURT
In multi-talker ASR, the input $\mathbf{X}\in\mathbb{R}^{T\times F}$ is an unsegmented mixture containing speech from multiple speakers (say, $K$ speakers), i.e., $\mathbf{X}=\sum_{n=1}^{N}\mathbf{x}_{n}$, where $\mathbf{x}_{n}$ is the $n$-th utterance ordered by start time, shifted and zero-padded to the length of $\mathbf{X}$. The desired output is $\mathbf{Y}=\bigcup_{n=1}^{N}\mathbf{y}_{n}$, where $\mathbf{y}_{n}$ is the reference corresponding to $\mathbf{x}_{n}$. If $N$ is small and fixed (e.g., $N=2$), a permutation-invariant training strategy may be used for this task. However, for general multi-talker scenarios, $N$ may be arbitrarily large.
Assuming at most two-speaker overlap, the heuristic error assignment training (HEAT) paradigm [20] is used to create channel-wise references $\mathbf{Y}_{1}$ and $\mathbf{Y}_{2}$ by assigning $\mathbf{y}_{n}$’s to the first available channel, in order of start time (Fig. 1). SURT estimates $\hat{\mathbf{Y}}=[\hat{\mathbf{Y}}_{1},\hat{\mathbf{Y}}_{2}]=f_{\text{surt}}(\mathbf{X})$ as follows. First, an unmixing module computes $\mathbf{H}_{1}$ and $\mathbf{H}_{2}$ as
@@
where $\bar{\mathbf{X}},\mathbf{M},\mathbbm{1}\in\mathbb{R}^{T\times D}$ (for latent dim. $D$), $\sigma$ is the sigmoid function, and $\ast$ is Hadamard product. $\mathbf{H}_{1}$ and $\mathbf{H}_{2}$ are fed into a transducer-based ASR, producing logits $\mathbf{Z}_{1}$ and $\mathbf{Z}_{2}$. Finally,
@@ IV-B Network architecture
We made several changes to the network architecture of SURT modules. For the masking network, we use dual-path LSTMs (DP-LSTMs) instead of 2-D convolutions in order to improve long range modeling capability for unmixing [89]. Second, in the encoder, we replace the DP-LSTMs (or DP-Transformers) with the recently proposed zipformer [90]. Zipformer contains multiple stacks of encoders running at different rates, with the middle ones more strongly down-sampled (by up to a factor of 4). This makes training more efficient as there are fewer frames to evaluate. Additionally, each block performs self-attention twice with shared attention weights, and a trainable bypass is introduced for each layer dimension. We also propose “branch tying” of the encoder outputs to jointly learn representation across all branches, i.e.,
@@ IV-C Training objective
Recall from Section III-A that the encoder and the prediction network generate hidden representations $\mathbf{f}_{1}^{T}$ and $\mathbf{g}_{1}^{U}$, respectively. In pruned transducer, these are first projected to $\mathbb{R}^{|\bar{\Sigma}|}$, and denoted as $\hat{\mathbf{f}}_{1}^{T}$ and $\hat{\mathbf{g}}_{1}^{U}$, respectively. A simple additive joiner is then used to compute trivial logits $\hat{\mathbf{z}}_{t,u}$ as
@@ 9                         $D_{\text{ovl}}$ = $D_{\text{ovl}}\cup\{-t\}$
25            offset = offset + ot $\mathcal{S}=\mathcal{S}\cup\{U_{i},\text{offset}\}$
26
@@ IV-D Mixture simulation
For SURT 2.0, we use sub-segments instead of full utterances as the source for mixture simulation, so that resulting mixtures are shorter while retaining multiple turns of conversation. These sub-segments are obtained using word-level alignment information, by breaking up the utterances at pauses longer than a threshold $\tau$. As an example, using $\tau=0.2$ for LibriSpeech resulted in sub-segments that were 2.8s on average. This allowed each training session to contain up to 9 turns of conversation while still being 36.5% shorter than the original training mixtures (Table I). Second, we learn histograms of pause/overlap distribution statistics from the target sessions, and sample from these distributions for mixing the segments. Such a strategy has been successfully applied to improve end-to-end neural diarization [94]. Our mixture simulation algorithm is described in Algorithm 1, and is similar to the conversation simulation algorithm from [94]. We assume that the input to the algorithm is the source segments $\mathcal{X}$, target sessions $\mathcal{M}$ (to learn statistics), maximum number of speakers $K$ in each mixture, and maximum duration $T$ of a speaker in a mixture. The algorithm returns the training mixtures $\mathcal{S}$. We optionally convolve $\mathcal{S}$ with real room impulse responses (RIRs) to train models for far-field reverberant conditions. The overall simulation workflow is shown in Fig. 3.
@@ V-B Evaluation
From Fig. 1, it is clear that SURT generates speaker-agnostic transcription with no strict correspondence between speaker and channel. This makes it impossible to evaluate SURT with the conventional word error rate (WER) metric used for single-speaker ASR systems. In [27] and [26], the authors independently proposed similar metrics that computed the minimum word error rate (WER) based on an optimal assignment of references to the channels, and this was termed optimal reference combination WER (ORC-WER) in [27]. In both studies, the authors remarked that computation of this metric was exponential in the number of reference utterances. Von Neumann et al. [98] proposed a polynomial-time implementation of ORC-WER using multi-dimensional Levenshtein distance, and released it through the meeteval [^4] https://github.com/fgnt/meeteval toolkit. They showed that ORC-WER is a lower bound on the popular concatenated minimum-permutation WER (cpWER) [6], and becomes equal to the cpWER when no speaker errors are present. We used their open-source implementation for evaluating our SURT models. The overall algorithm for ORC-WER computation is given in Algorithm 2. We assume that the references $\mathcal{R}_{n}$ are ordered in increasing order of start times. In the remainder of this paper, we will use the abbreviation WER to actually mean ORC-WER, unless explicitly mentioned. For LibriCSS, we report WERs on the officially provided (approximately 1 min. long) “segments,” following prior work on continuous input evaluation [57, 60, 26]. For AMI and ICSI, we use utterance-group based evaluation similar to [54, 99]. This results in dev/test segments of duration 7.0s/8.4s and 2.6s/2.9s for AMI and ICSI, respectively.
Input: Ref.: $\mathcal{R}_{1},\ldots,\mathcal{R}_{N}$; hyp.: $\mathcal{H}_{1},\ldots,\mathcal{H}_{C}$
Output: $W_{\text{orc}}$
$\mathcal{R}\leftarrow\mathcal{R}_{1}\diamond\ldots\diamond\mathcal{R}_{N}$   // Insert channel change token
$\xi\leftarrow\{\}$   // Memoize costs to avoid recomputing
1 Function levenshtein($R$, $H$, $i=\phi$):
2       $K=(R,H,i)$ if $K$ in $\xi$ then
3             return $\xi[K]$
4      if $i\neq\phi$ and $\mathcal{R}[R-1]==\diamond$ then
              // Reached a channel change token
5             $\xi[K]\leftarrow$ levenshtein($R-1$,$H$,$\phi$) return $\xi[K]$
6      if $i=\phi$ then
             $C\leftarrow\emptyset$   // Find minimum over all hypotheses
7             for $i$ in $|\mathcal{H}|$) do
8                   $C[i]\leftarrow$ levenshtein($R$,$H$,$i$)
9            $\xi[K]\leftarrow\min(C)$ return $\xi[K]$
10      else
              // Regular Levenshtein update
11             if $R=0$ then
12                   $c\leftarrow\sum_{h\in H}h$
13             else if $H[i]=0$ then
14                   $c\leftarrow 1$ + levenshtein($R-1$,$H$,$i$)
15             else if $\mathcal{R}[R-1]=\mathcal{H}[i][H[i]-1]$ then
16                   $H[i]\leftarrow H[i]-1$ $c\leftarrow$ levenshtein($R-1$,$H$,$i$) $H[i]\leftarrow H[i]+1$
17             else
18                   $ins\leftarrow$ levenshtein($R-1$,$H$,$i$) $H[i]\leftarrow H[i]-1$ $del\leftarrow$ levenshtein($R$,$H$,$i$) $sub\leftarrow$ levenshtein($R-1$,$H$,$i$) $H[i]\leftarrow H[i]+1$ $c\leftarrow 1+\min(ins,del,sub)$
19             $\xi[K]\leftarrow c$ return $\xi[K]$
20
return $|\mathcal{R}|$,$\{|h|:h\in\mathcal{H}\}$)
Algorithm 2 ORC-WER computation
@@ V-C Implementation details
Network architecture. We experimented with two variants of SURT — base and large. The base model contains four 256-dim DP-LSTM layers trained with chunk width randomization [26] as the masking network. The encoder consists of 5 zipformer blocks with 2 self-attention layers per block. Each block consists of a 192-dim attention distributed across 8 heads, and a 768-dim feed-forward layer. Downsampling factors of (1,2,4,8,2) were used in the zipformer blocks. The prediction network contains a single 512-dim Conv1D layer. The large model contains 6 layers in the masking network, and (2,4,3,2,4) self-attention layers in the 5 zipformer blocks.
@@
Training data. For LibriCSS experiments, we first created anechoic mixtures, LSMix-clean, using each speed-perturbed LibriSpeech train sub-segment ($\tau=0.2$; cf. § IV-D) once, resulting in approx. 2200h of training data. A reverberated copy of LSMix, named LSMix-reverb, was generated by convolving LSMix with real RIRs collected from the REVERB [5] dataset. We also added isotropic noises from the REVERB data, and perturbed the loudness between -20 dB and -25 dB using the pyloudnorm tool [100]. We will refer to the combination of LSMix-clean and LSMix-reverb as LSMix-full hereafter. Ablation experiments were conducted on the anechoic LibriCSS by training SURT on LSMix-clean. For these experiments, we used on-the-fly noise augmentation using noises from the MUSAN corpus [101]. For final evaluation, we trained SURT on LSMix-full ($\sim$4400h), so that the same model can be used on both anechoic and replayed LibriCSS. All models were trained using on-the-fly SpecAugment [102]. We found it beneficial to use high overlap mixtures in the warm-up stage of SURT training to encourage better mask estimation [103]. For pre-training on single-speaker data, we used LibriSpeech train set, optionally convolved with synthetic RIRs (for the final evaluation). This pre-training was done for 10 epochs.
@@
Since AMI and ICSI have similar characteristics, we trained a combined SURT model for them using synthetic mixtures created from close-talk utterances, again using sub-segments obtained from forced alignments ($\tau=0.5$). We refer to these mixtures as AIMix-clean, their reverberant copy as AIMix-reverb, and the combination as AIMix-full. The models were subsequently adapted by combining the real train sessions from all microphone settings.
@@
Hyper-parameters. The auxiliary loss scales, $\lambda_{\text{ctc}}$ and $\lambda_{\text{mask}}$, were set to 0.2 each. $\mathcal{L}_{\text{mask}}$ was not used during adaptation since ground-truth separated audio may not be available for real data. We trained the models with the ScaledAdam optimizer following the standard zipformer-transducer recipes in icefall. This a variant of Adam where each parameter’s update is scaled proportional to the norm of that parameter. The learning rate was warmed up to 0.004 for 5000 iterations, and decayed exponentially thereafter. All models were trained for 30 epochs using 4 GPUs [^5] We used either Titan RTX (with batch size 500s) or V100 (with batch size 650s) depending on availability on our compute cluster..

### v2

@@ I Introduction
For the case of streaming multi-talker ASR, two methods have been proposed to extend neural transducers [20] for the task. These methods are beneficial since neural transducers (using RNNs or transformers) have become the standard modeling technique for on-device speech recognition [21, 22, 23]. The first method, exemplified by token-level SOT (t-SOT) [24], uses a conventional single-branch model similar to single-talker ASR, whereas the second method relies on a two-branch modeling strategy, as exemplified by streaming unmixing and recognition transducer (SURT) [25] and multi-turn RNNT (MT-RNNT) [26]. In this paper, we focus on the latter class of models, and refer to them as SURT, but the same ideas should also be applicable to MT-RNNT. The SURT model separates overlapping speech into multiple simultaneous branches (or channels), each of which is transcribed by a shared transducer. It has been extended to handle long-form multi-turn recordings [27, 28], and to jointly perform speaker identification [29], endpointing [30], and segmentation [31].
@@ III-A Speech recognition with neural transducers
In conventional single-talker ASR, audio features for a segmented utterance $\mathbf{X}\in\mathbb{R}^{T\times F}$ (where $T$ and $F$ denote the number of time frames and the input feature dimension, respectively) are provided as input to the system, and we are required to predict the transcript $\mathbf{y}=(y_{1},\ldots,y_{U})$, where $y_{u}\in\mathcal{V}$ denotes output units such as graphemes or word-pieces, and $U$ is the length of the label sequence. For the case of discriminative training, this requires computing the conditional likelihood $P(\mathbf{y}|\mathbf{X})$ (or its log for numerical stability). For inference, we search for $\hat{\mathbf{y}}=\text{arg}\max_{\mathbf{y}}P(\mathbf{y}|\mathbf{X})$, often in a constrained search space using greedy or beam search. Transducers achieve this by marginalizing over the set of all alignments $\mathbf{a}\in\bar{\mathcal{V}}^{T+U}$, where $\bar{\mathcal{V}}=\mathcal{V}\cup\{\phi\}$ and $\phi$ is called the blank label. Formally,
@@
Transducers parameterize $P(\mathbf{a}|\mathbf{X})$ with an encoder, a prediction network, and a joiner (see “recognition” component in Fig. 2a). The encoder maps $\mathbf{x}$ into hidden representations $\mathbf{f}_{1}^{T}$, while the prediction network maps $\mathbf{y}$ into $\mathbf{g}_{1}^{U}$. The joiner combines the outputs from the encoder and the prediction network to compute logits $\mathbf{z}_{t,u}$ which are fed to a softmax function to produce a posterior distribution over $\bar{\mathcal{V}}$. Under assumptions of full context encoder, we can expand (1) as
@@
where $u(t)\in\{1,\ldots,U\}$ denotes the index in the label sequence at time $t$. We will denote the log of this expression as $\mathcal{L}_{\text{rnnt}}(\mathbf{y},\mathbf{z})$ (or simply $\mathcal{L}_{\text{rnnt}}$) for the remainder of this paper.
@@ III-B Multi-talker ASR with SURT
In multi-talker ASR, the input $\mathbf{X}\in\mathbb{R}^{T\times F}$ is an unsegmented mixture containing $N$ utterances from $K$ speakers, i.e., $\mathbf{X}=\sum_{n=1}^{N}\mathbf{x}_{n}$, where $\mathbf{x}_{n}$ is the $n$-th utterance ordered by start time, shifted and zero-padded to the length of $\mathbf{X}$. The desired output is $\mathbf{Y}=\{\mathbf{y}_{n}:1\leq n\leq N\}$, where $\mathbf{y}_{n}$ is the reference corresponding to $\mathbf{x}_{n}$. If $N$ is small and fixed (e.g., $N=2$), a permutation-invariant training strategy may be used for this task. However, for general multi-talker scenarios, $N$ may be arbitrarily large.
Assuming at most two-speaker overlap, the heuristic error assignment training (HEAT) paradigm [25] is used to create channel-wise references $\mathbf{Y}_{1}$ and $\mathbf{Y}_{2}$ by assigning $\mathbf{y}_{n}$’s to the first available channel, in order of start time (Fig. 1). Formally, if $\zeta:n\rightarrow\{1,2\}$ maps utterances to channels, $\theta_{n}^{\mathrm{st}}$ and $\theta_{n}^{\mathrm{en}}$ denote the start and end times for utterance $n$, and $\theta_{n}^{\mathrm{st}}$ is monotonically increasing $\forall n$, we have
$\zeta(n)=\begin{cases}1,&\text{if}~{}~{}\theta_{n}^{\mathrm{st}}\geq\max_{i\in\zeta^{-1}(1)}\theta_{i}^{\mathrm{en}}\\ 2,&\text{otherwise},\end{cases}$ (4)
and $\zeta(n)$’s are assigned sequentially. SURT estimates $\hat{\mathbf{Y}}=[\hat{\mathbf{Y}}_{1},\hat{\mathbf{Y}}_{2}]=f_{\text{surt}}(\mathbf{X})$ as follows. First, an unmixing module computes $\mathbf{H}_{1}$ and $\mathbf{H}_{2}$ as
@@
where $\bar{\mathbf{X}},\mathbf{M},\mathbbm{1}\in\mathbb{R}^{T\times D}$ (for latent dim. $D$) is a matrix of ones, $\sigma$ is the sigmoid function, and $\ast$ is Hadamard product. $\mathbf{H}_{1}$ and $\mathbf{H}_{2}$ are fed into a transducer-based ASR, producing logits $\mathbf{Z}_{1}$ and $\mathbf{Z}_{2}$. Finally,
@@ IV-B Network architecture
Figure 3: Illustration of the zipformer encoder architecture. The encoder contains multiple “blocks” running at different frame rates (left). Each block contains several “layers” (middle), and each layer performs self-attention twice with shared attention weights (right).
We made several changes to the network architecture of SURT modules. For the masking network, we use dual-path LSTMs (DP-LSTMs) instead of 2-D convolutions in order to improve long range modeling capability for unmixing [90]. Second, in the encoder, we replace the DP-LSTMs (or DP-Transformers) with the recently proposed zipformer [91]. As shown in Fig. 3, the zipformer encoder contains multiple encoder blocks running at different rates, with the middle ones more strongly down-sampled (by up to a factor of 8). This makes training more efficient as there are fewer frames to evaluate. Each block may contain one or more encoder “layers” operating at the same frame rate. Each layer performs self-attention twice with shared attention weights, and a trainable bypass is introduced for each layer dimension. We also propose “branch tying” of the encoder outputs to jointly learn representation across all branches, i.e.,
@@ IV-C Training objective
Recall from Section III-A that the encoder and the prediction network generate hidden representations $\mathbf{f}_{1}^{T}$ and $\mathbf{g}_{1}^{U}$, respectively. In pruned transducer, these are first projected to $\mathbb{R}^{|\bar{\mathcal{V}}|}$, and denoted as $\hat{\mathbf{f}}_{1}^{T}$ and $\hat{\mathbf{g}}_{1}^{U}$, respectively. A simple additive joiner is then used to compute trivial logits $\hat{\mathbf{z}}_{t,u}$ as
@@ 9                         $D_{\text{ovl}}$ = $D_{\text{ovl}}\cup\{-t\}$
       $\mathcal{S}_{\mathrm{cur}}\leftarrow\phi$  // initialize empty mixture
@@
25            offset = offset + ot $\mathcal{S}_{\mathrm{cur}}=\mathcal{S}_{\mathrm{cur}}\cup\{U_{i},\text{offset}\}$
26      $\mathcal{S}=\mathcal{S}\cup\mathcal{S}_{\mathrm{cur}}$
@@ IV-D Mixture simulation
For SURT 2.0, we use sub-segments instead of full utterances as the source for mixture simulation, so that resulting mixtures are shorter while retaining multiple turns of conversation. These sub-segments are obtained using word-level alignment information, by breaking up the utterances at pauses longer than a threshold $\tau$. As an example, using $\tau=0.2$ for LibriSpeech resulted in sub-segments that were 2.8s on average. This allowed each training session to contain up to 9 turns of conversation while still being 36.5% shorter than the original training mixtures (Table I). Second, we learn histograms of pause/overlap distribution statistics from the target sessions, and sample from these distributions for mixing the segments. Such a strategy has been successfully applied to improve end-to-end neural diarization [95]. Our mixture simulation algorithm is described in Algorithm 1, and is similar to the conversation simulation algorithm from [95]. We assume that the input to the algorithm is the source segments $\mathcal{X}$, target sessions $\mathcal{M}$ (to learn statistics), maximum number of speakers $K$ in each mixture, and maximum duration $T$ of a speaker in a mixture. The algorithm returns the training mixtures $\mathcal{S}$. For training, the input speech mixture $\mathbf{X}$ is obtained by digitally adding the utterances in $\mathcal{S}$ with the specified offsets. The utterance-wise labels $\mathbf{y}_{n}$, along with $\theta_{n}^{\mathrm{st}}$ and $\theta_{n}^{\mathrm{en}}$ (uniquely determined by the offset and duration of utterance $n$) are used to obtain the reference transcripts for both channels using equation (4). We optionally convolve $\mathcal{S}$ with real room impulse responses (RIRs) to train models for far-field reverberant conditions. The overall simulation workflow is shown in Fig. 4.
@@ V-B Evaluation
Fig. 1 shows that SURT generates speaker-agnostic transcription with no strict correspondence between speaker and channel. This makes it impossible to evaluate SURT with the conventional word error rate (WER) metric used for single-speaker ASR systems. [27] and [28] independently proposed similar metrics that computed the minimum word error rate (WER) based on an optimal assignment of references to the channels, and this was termed optimal reference combination WER (ORC-WER) in [28]. In both studies, the authors remarked that computation of this metric was exponential in the number of reference utterances. Von Neumann et al. [99] proposed a polynomial-time implementation of ORC-WER using multi-dimensional Levenshtein distance, and released it through the meeteval [^4] https://github.com/fgnt/meeteval toolkit. They showed that ORC-WER is a lower bound on the popular concatenated minimum-permutation WER (cpWER) [6], and becomes equal to the cpWER when no speaker errors are present. We used their open-source implementation for evaluating our SURT models. In the remainder of this paper, we will use the abbreviation WER to actually mean ORC-WER, unless explicitly mentioned. For LibriCSS, we report WERs on the officially provided (approximately 1 min. long) “segments,” following prior work on continuous input evaluation [57, 60, 27]. For AMI and ICSI, we use utterance-group based evaluation similar to [54, 100]. This results in dev/test segments of duration 7.0s/8.4s and 2.6s/2.9s for AMI and ICSI, respectively.
@@ V-C Implementation details
Network architecture. We experimented with two variants of SURT — base and large. The base model contains four 256-dim DP-LSTM layers trained with chunk width randomization [27] as the masking network. The encoder consists of 5 zipformer blocks with 2 self-attention layers per block. Each block consists of a 192-dim attention distributed across 8 heads, and a 768-dim feed-forward layer. Downsampling factors of (1,2,4,8,2) were used in the zipformer blocks. The prediction network contains a single 512-dim Conv1D layer. The large model contains 6 layers in the masking network, and (2,4,3,2,4) self-attention layers in the 5 zipformer blocks. The chunk size for the intra-LSTM and the Zipformer is set to 32 frames, resulting in a modeling latency of 320 ms.
@@
Training data. For LibriCSS experiments, we first created anechoic mixtures, LSMix-clean, using each speed-perturbed LibriSpeech train sub-segment ($\tau=0.2$; cf. § IV-D) once, resulting in approx. 2200h of training data. For this, we set $\mathcal{M}$ as the LibriCSS dev set (excluding the 0L and OV10 sessions), $K=3$, and $T=15$ in Algorithm 1. A reverberated copy of LSMix, named LSMix-reverb, was generated by convolving LSMix with real RIRs collected from the REVERB [5] dataset. We also added isotropic noises from the REVERB data, and perturbed the loudness between -20 dB and -25 dB using the pyloudnorm tool [101]. We will refer to the combination of LSMix-clean and LSMix-reverb as LSMix-full hereafter. Ablation experiments were conducted on the anechoic LibriCSS by training SURT on LSMix-clean. For these experiments, we used on-the-fly noise augmentation using noises from the MUSAN corpus [102]. For final evaluation, we trained SURT on LSMix-full ($\sim$4400h), so that the same model can be used on both anechoic and replayed LibriCSS. All models were trained using on-the-fly SpecAugment [103]. We found it beneficial to use high overlap mixtures in the warm-up stage of SURT training to encourage better mask estimation [104]. For pre-training on single-speaker data, we used LibriSpeech train set, optionally convolved with synthetic RIRs (for the final evaluation). This pre-training was done for 10 epochs.
@@
Since AMI and ICSI have similar characteristics, we trained a combined SURT model for them using synthetic mixtures created from close-talk utterances, again using sub-segments obtained from forced alignments ($\tau=0.5$). These mixtures were obtained by setting $D_{=\text{spk}}$, $D_{\neq\text{spk}}$, $D_{\text{ovl}}$, and $P_{\text{ovl}}$ as 0.5, 0.5, 1.0, and 0.8, respectively, in Algorithm 1. $K$ and $T$ were set to 3 and 15s, respectively, similar to LSMix-clean simulation. We refer to these mixtures as AIMix-clean, their reverberant copy as AIMix-reverb, and the combination as AIMix-full. The models were subsequently adapted by combining the real train sessions from all microphone settings.
@@
Hyper-parameters. The auxiliary loss scales, $\lambda_{\text{ctc}}$ and $\lambda_{\text{mask}}$, were set to 0.2 each. $\mathcal{L}_{\text{mask}}$ was not used during adaptation since ground-truth separated audio may not be available for real data. We trained the models with the ScaledAdam optimizer following the standard zipformer-transducer recipes in icefall. This is a variant of Adam where each parameter’s update is scaled proportional to the norm of that parameter. The learning rate was warmed up to 0.004 for 5000 iterations, and decayed exponentially thereafter. All models were trained for 30 epochs using 4 GPUs [^5] We used either Titan RTX (with batch size 500s) or V100 (with batch size 650s) depending on availability on our compute cluster..